4C2 is equal to _____________


1
Expert's answer
2021-06-01T17:47:04-0400

We know that,


"^nC_r=\dfrac{n!}{r!\cdot(n-r)!}"

So,



"^4C_2=\dfrac{4!}{2!\cdot (4-2)!}=\dfrac{4!}{2!\times2!~}=\dfrac{24}{4}=6"

Hence,


"\boxed{^4C_2=6}"